# 3DCoMPaT++: An improved Large-scale 3D Vision Dataset for Compositional Recognition

### v2

@@ Abstract
In this work, we present 3DCoMPaT++, a multimodal 2D/3D dataset with 160 million rendered views of more than 10 million stylized 3D shapes carefully annotated at the part-instance level, alongside matching RGB point clouds, 3D textured meshes, depth maps, and segmentation masks. 3DCoMPaT++ covers 41 shape categories, 275 fine-grained part categories, and 293 fine-grained material classes that can be compositionally applied to parts of 3D objects. We render a subset of one million stylized shapes from four equally spaced views as well as four randomized views, leading to a total of 160 million renderings. Parts are segmented at the instance level, with coarse-grained and fine-grained semantic levels. We introduce a new task, called Grounded CoMPaT Recognition (GCR), to collectively recognize and ground compositions of materials on parts of 3D objects. Additionally, we report the outcomes of a data challenge organized at the CVPR conference, showcasing the winning method’s utilization of a modified PointNet++ model trained on 6D inputs, and exploring alternative techniques for GCR enhancement. We hope our work will help ease future research on compositional 3D Vision. The dataset and code have been made publicly available at https://3dcompat-dataset.org/v2/.
@@ 1 Introduction
Dataset. To create our dataset, we start with 10K unique geometries which we segment at a fine-grained part level into a total of 275 segmented parts (leading to 9.98 average part instances per shape). For each part of each shape, human experts determine a list of compatible/applicable materials. Then, we generate a stylized model by sampling over the compatible materials of each part with a limit of 1000 styles per shape, leading to 10 million stylized shapes.
@@
(c) Texture coordinates, textures, and materials. A key focus of our work lies in the stylization of 3D shapes with appropriate texture coordinates, textures, and materials. To achieve a superior level of quality when rendering numerous material compositions on each shape, our models are equipped with human-verified texture coordinates and part-wise material compatibility information. While previous attempts have been made to enhance a subset of ShapeNet with part-wise material information [13], it falls short in comparison to our work in terms of the number of shapes (3080 vs. 10000), shape classes (3 vs 41), and materials (6 vs. 13, and 293 fine-grained annotated classes).
@@ 2 Related work
[…]
Figure 3: Statistics of the number of edges, vertices, and faces (left), Density plots depicting the distribution of vertex counts, edge counts, and face counts across 3D shapes extracted from both the 3DCoMPaT++ and ShapeNet datasets (right). Results highlight significantly higher numbers of vertices, edges, and faces in 3DCoMPaT++ shapes compared to ShapeNet. Note that all annotated shapes in PartNet [3] originate from ShapeNet.
@@
Part-understanding. In an attempt to bridge this gap, ShapeNet-Part [11] was first proposed as an extension to ShapeNet [1] with part-level annotations. It contains 16 shape categories and 31K 3D shapes, but part annotations are only provided at a coarse-grained semantic level, and are extracted using a deep active learning framework instead of human annotation. PhotoShape [15] is one of the earliest efforts in gathering 3D shapes with high-quality textures. It contains 5.8K 3D shapes from 29 shape categories, and proposes to transfer materials properties regressed from real images to untextured 3D shapes. PartNet [3] was built as a large-scale dataset of 3D shapes annotated with fine-grained, instance-level, and hierarchical part segmentations. PartNet is also created on top of ShapeNet [1] and contains 26K 3D shapes from 24 shape categories. PartNet is a valuable resource for advancing research in 3D shape analysis and understanding. Our work stands apart from PartNet in three main significant ways:
@@
  • (c) We use a human verification process to ensure the compatibility of sampled materials with each shape/part.
High-resolution datasets. In an effort to provide high-quality and realistic shapes and textures, OmniObject3D [4] and ABO [5] datasets introduced 3D assets with rich, high-quality textures. Google Scanned Objects [16] is a scanned dataset of reconstructed 3D objects with high-quality textures and geometries, and contains 1K 3D shapes from 17 diverse categories of small objects. OmniObject3D [4] is a scanned dataset of 3D objects with high-quality textures, and contains 6K 3D shapes from 190 shape categories based on ImageNet [19] and LVIS [20]. ABO [5] is a dataset of 3D objects with high-quality textures and geometries, and contains 8K 3D shapes from 63 shape categories based on product catalogs extracted from Amazon.
@@
Mesh resolutions. In Figure 3, we compare model resolution statistics for 3D CAD models from 3DCoMPaT++ and ShapeNet [1]. This comparison is important because ShapeNet [1] serves as the CAD model data source for PartNet [3] and ShapeNet-Part [11], which are two of the most prominent datasets for 3D part understanding. We provide density plots over vertex counts, edges counts, and faces counts for 3D CAD models from both datasets. We also provide the median values of these metrics for both datasets. We show that 3DCoMPaT++ exhibits shapes with significantly higher average numbers of vertices, edges, and faces when compared to ShapeNet and PartNet. While polygon count is not a perfect proxy for shape visual quality and realism, it is a useful metric for comparing the relative complexity of meshes in each dataset. This quantitative assessment underscores the richness of geometries within 3DCoMPaT++, making it a valuable resource for advancing research in 3D shape analysis and understanding.
Figure 5: Rendering of randomly sampled shapes from 3DCoMPaT++. The dataset comprises a rich collection of stylized 3D shapes annotated at the part-instance level. These renderings demonstrate the varying shapes, styles, and materials that are captured, enabling comprehensive exploration and analysis of compositional 3D vision tasks. Shapes are consistently aligned across classes and orientations are consistent for all 3D models. In the left circle, we illustrate the untextured 3D geometries we start from as a reference.
Figure 6: Canonical (left) and random (right) views rendered for a single stylized 3D shape. Random viewpoints are sampled, while canonical viewpoints are equally spaced around the shape. In Figure 17 of the appendix, we provide additional examples of canonical and random views.
3 3DCoMPaT++
The 3DCoMPaT++ dataset is based on a collection of artist-designed 3D CAD models collected and annotated in collaboration with an industry partner. It contains 10K geometries annotated and segmented at a fine-grained part-instance level, with material compatibility information for each annotated part. For each shape, 8 rendered views are provided from canonical and random viewpoints (see Figure 6). For each rendered view, depth maps, part maps, and material maps are rendered (see Figures 7 and 18).
@@ 3.1 Dataset
Figure 8: Distribution of part occurrences for fine/coarse levels. We plot the sorted number of occurrences of each part (logscale, top). We also plot the average number of unique parts per object for fine and coarse semantic levels, across all shape categories (bottom).
@@
3D Data. Alongside each stylized shape, we provide a part-segmented textured 3D mesh, an RGB pointcloud, and point-wise part and material annotations. All part segmentation information is provided in coarse-grained and fine-grained semantic levels. RGB pointclouds can be resampled at any resolution starting from the available textured 3D meshes. In Figure 1, we illustrate the 3D data provided for a single stylized shape.
@@
3.2 Data collection pipeline
@@
  • Rendering. We render each shape from multiple viewpoints with matching masks, depth maps and pointcloud data, as detailed in Section 3.4.
@@
3.3 Coarse/Fine-grained semantics
@@
We plot in Figure 8 the sorted number of occurrences of each part (top) for fine and coarse semantic levels. We also compare the average number of unique parts per object (bottom). In the coarse-grained semantic level, parts occurrences are concentrated on a smaller number of parts, while the distribution for the fine-grained level is clearly long-tailed. The average number of parts per object is also equalized across shape categories in the coarse-grained level, while some shape categories present a significantly higher number of parts per object in the fine-grained level like cars and bicycles. Overall, the coarse-grained semantic level provides a more balanced distribution of parts across shape categories, while the fine-grained semantic level provides a more detailed description of the composition of shapes.
@@
ceramic
fabric
glass
granite
leather
marble
metal
@@
plant
plastic
rubber
soil
wax
wood
Total
@@ 3.4 Rendering
All images are rendered in a 256x256 resolution, and 2D images are encoded in the PNG format. Depth maps are stored in the OpenEXR format to accommodate absolute distances to the image plane, which are represented using floating-point values.
Viewpoints. Each stylized shape is rendered from multiple perspectives: 4 canonical viewpoints and 4 random viewpoints. We first translate each shape above the $z=0$ plane. Camera viewpoints are defined in spherical coordinates $(\phi,\theta)$ where the origin is set to the center of the shape’s bounding box, which we note $\mathbf{o}_{c}$. The camera is rotated around $\mathbf{o}_{c}$ by $\phi$ and $\theta$. Canonical viewpoints are distributed evenly around the shape with a fixed elevation $\theta$. We set the base spherical angle $\phi$ to 40 degrees and then increment it by 90 degrees for each of the four viewpoints, while keeping $\theta$ fixed at 0 degrees. Random viewpoints are sampled uniformly from an upper hemisphere above the plane. We randomly sample $\phi$ from the range $[0,2\pi]$ and $\theta$ from the range $[-\frac{1}{3}\pi,\frac{1}{3}\pi]$. Using the obtained $\phi$ and $\theta$ angles, we define the position and orientation of the camera. The camera’s initial position denoted as $\mathbf{c_{0}}$ is rotated around $\mathbf{o_{c}}$. The orientation of the camera is then adjusted to ensure that the image plane is centered on $\mathbf{o_{c}}$. Extrinsic and intrisic camera parameters are recorded for each view and are provided alongside the rendered images.
The sampling procedure of camera parameters is detailed in Algorithm 1.
Algorithm 1  Sampling camera parameters
Input:
Output:
[…]
3.5 Toolbox
@@
4 Experiments
4.1 Classification and Segmentation
@@
Part segmentation. We conduct 3D part segmentation experiments on pointclouds and 2D renders to assess the difficulty of this task on our dataset. We provide results for both fine-grained and coarse-grained 3D part segmentation in Table III. We report pointwise accuracy (shape-agnostic) and mIOU for each model. For mIOU, we consider the shape-informed version of the metric where we restrict the set of predicted parts to the parts that are present in the ground-truth shape category, and the shape-agnostic version where all possible parts are considered. We also report results with and without using a shape prior during training and inference for PCT [27], PointNet++ [28] and CurveNet [26]. We note that getting accurate part segmentations without RGB information is challenging but remains possible. Without using a shape prior, CurveNet [26] reaches a shape-agnostic mIOU of 53.09% on fine-grained part segmentation. In this setting, the model has to perform the challenging task of point-wise part classification from a set of 275 possible parts.
@@
Overall, a large gap exists (around 30 accuracy points across models) between shape-informed and shape-agnostic mIOU, highlighting the difficulty of the task over the full space of possible parts. The task of coarse-grained part segmentation is easier, as the model only has to perform part classification from a set of 43 possible parts. In this setting, CurveNet [26] reaches a shape-agnostic mIOU of 76.32%. For 2D fine and coarse part segmentation, we report results for SegFormer [29] in Table IV, alongside material segmentation results. We obtain a mIOU of 52.24% for fine-grained part segmentation, 73.35% for coarse-grained part segmentation, and 82.45% for material segmentation.
4.2 Grounded Compositional Recognition (GCR)

### v3

@@ Abstract
In this work, we present 3DCoMPaT++, a multimodal 2D/3D dataset with 160 million rendered views of more than 10 million stylized 3D shapes carefully annotated at the part-instance level, alongside matching RGB point clouds, 3D textured meshes, depth maps, and segmentation masks. 3DCoMPaT++ covers 42 shape categories, 275 fine-grained part categories, and 293 fine-grained material classes that can be compositionally applied to parts of 3D objects. We render a subset of one million stylized shapes from four equally spaced views as well as four randomized views, leading to a total of 160 million renderings. Parts are segmented at the instance level, with coarse-grained and fine-grained semantic levels. We introduce a new task, called Grounded CoMPaT Recognition (GCR), to collectively recognize and ground compositions of materials on parts of 3D objects. Additionally, we report the outcomes of a data challenge organized at the CVPR conference, showcasing the winning method’s utilization of a modified PointNet++ model trained on 6D inputs, and exploring alternative techniques for GCR enhancement. We hope our work will help ease future research on compositional 3D Vision. The dataset and code have been made publicly available at https://3dcompat-dataset.org/v2/.
@@ I Introduction
Dataset. To create our dataset, we start with 10K unique geometries which we segment at a fine-grained part level into a total of 275 segmented parts (leading to 9.86 average part instances per shape). For each part of each shape, human experts determine a list of compatible/applicable materials. Then, we generate a stylized model by sampling over the compatible materials of each part with a limit of 1000 styles per shape, leading to 10 million stylized shapes.
@@
(c) Texture coordinates, textures, and materials. A key focus of our work lies in the stylization of 3D shapes with appropriate texture coordinates, textures, and materials. To achieve a superior level of quality when rendering numerous material compositions on each shape, our models are equipped with human-verified texture coordinates and part-wise material compatibility information. While previous attempts have been made to enhance a subset of ShapeNet with part-wise material information [13], it falls short in comparison to our work in terms of the number of shapes (3080 vs. 10000), shape classes (3 vs 42), and materials (6 vs. 13, and 293 fine-grained annotated classes).
@@ II Related work
Figure 3: Comparison with PartNet. Part instances per shape distributions compared to PartNet [3] (left), Density plots depicting the distribution of vertex counts, edge counts, and face counts across 3D shapes extracted from both the 3DCoMPaT++ and ShapeNet datasets (right). We show significantly higher numbers of vertices, edges, and faces in 3DCoMPaT++ shapes compared to ShapeNet. All annotated shapes in PartNet originate from ShapeNet.
@@
Part-understanding. In an attempt to bridge this gap, ShapeNet-Part [11] was first proposed as an extension to ShapeNet [1] with part-level annotations. It contains 16 shape categories and 31K 3D shapes, but part annotations are only provided at a coarse-grained semantic level, and are extracted using a deep active learning framework instead of human annotation. PhotoShape [15] is one of the earliest efforts in gathering 3D shapes with high-quality textures. It contains 5.8K 3D shapes from 29 shape categories, and proposes to transfer material properties regressed from real images to untextured 3D shapes. PartNet [3] was built as a large-scale dataset of 3D shapes annotated with fine-grained, instance-level, and hierarchical part segmentations. PartNet is also created on top of ShapeNet [1] and contains 26K 3D shapes from 24 shape categories. PartNet is a valuable resource for advancing research in 3D shape analysis and understanding. Our work stands apart from PartNet in three main ways:
@@
  • (c) We use a human verification process to ensure the compatibility of sampled materials with each part of each shape.
High-resolution datasets. In an effort to provide high-quality and realistic shapes and textures, OmniObject3D [4] and ABO [5] datasets introduced 3D assets with rich, high-quality textures. ABO [5] also provides unlabelled part instance segments for a subset of 3,4K shapes, in the form of unnamed connected shape pieces. HAL3D [18] builds on ABO using an active learning pipeline to provide semantic part instance names for ABO, but did not release the dataset to the public. Google Scanned Objects [16] is a scanned dataset of reconstructed 3D objects with high-quality textures and geometries, and contains 1K 3D shapes from 17 diverse categories of small objects. OmniObject3D [4] is a scanned dataset of 3D objects with high-quality textures, and contains 6K 3D shapes from 190 shape categories based on ImageNet [20] and LVIS [21]. ABO [5] is a dataset of 3D objects with high-quality textures and geometries, and contains 8K 3D shapes from 63 shape categories based on product catalogs extracted from Amazon.
@@
Mesh resolutions. In Figure 3, we compare part instances per shape distributions between 3DCoMPaT++ and PartNet [3] (left), and model resolution statistics for 3D CAD models from 3DCoMPaT++ and ShapeNet [1] (right). This comparison is important because ShapeNet [1] serves as the CAD model data source for PartNet [3] and ShapeNet-Part [11], which are two of the most prominent datasets for 3D part understanding. We provide density plots over vertex counts, edges counts, and faces counts for 3D CAD models from both datasets. We show that 3DCoMPaT++ exhibits both higher numbers of part instances per shape compared to PartNet and significantly higher average numbers of vertices, edges, and faces when compared to ShapeNet. While polygon count is not a perfect proxy for shape visual quality and realism, it is a useful metric for comparing the relative complexity of meshes in each dataset. This quantitative assessment underscores the richness of annotations and geometries within 3DCoMPaT++, making it a valuable resource for advancing research in 3D shape analysis and understanding.
Figure 5: Rendering of randomly sampled shapes from 3DCoMPaT++. The dataset comprises a rich collection of stylized 3D shapes annotated at the part-instance level. These renderings demonstrate the varying shapes, styles, and materials that are captured, enabling comprehensive exploration and analysis of compositional 3D vision tasks. Shapes are consistently aligned across classes and orientations are consistent for all 3D models. In the left circle, we illustrate the untextured 3D geometries we start from as a reference. We provide additional reference shapes from all 42 shape categories in Figure 17 of the appendix.
Figure 6: Canonical and random views. Canonical (left) and random (right) views rendered for a single stylized 3D shape. Random viewpoints are sampled, while canonical viewpoints are equally spaced around the shape. In Figure 20 of the appendix, we provide additional examples of canonical and random views.
III 3DCoMPaT++
The 3DCoMPaT++ dataset is based on a collection of artist-designed 3D CAD models collected and annotated in collaboration with an industry partner. It contains 10K geometries annotated and segmented at a fine-grained part-instance level, with material compatibility information for each annotated part. For each shape, 8 rendered views are provided from canonical and random viewpoints (see Figure 6). For each rendered view, depth maps, part maps, and material maps are rendered (see Figures 7 and 21).
@@ III-A Dataset
Figure 8: Distribution of part occurrences for fine/coarse levels. We plot the average number of unique parts per object for fine and coarse semantic levels, across all shape categories (top). We also plot the sorted number of occurrences of each part (logscale, bottom).
@@
3D Data. Alongside each stylized shape, we provide a part-segmented textured 3D mesh, an RGB pointcloud, and point-wise and triangle-wise part and material annotations. All part segmentation information is provided in coarse-grained and fine-grained semantic levels. RGB pointclouds can be resampled at any resolution starting from the available textured 3D meshes. In Figure 1, we illustrate the 3D data provided for a single stylized shape.
@@
III-B Data collection pipeline
@@
  • Rendering. We render each shape from multiple viewpoints with matching masks, depth maps and pointcloud data, as detailed in Section III-D.
@@
III-C Coarse/Fine-grained semantics
@@
We compare the average number of unique parts per object (top) for fine and coarse semantic levels across all shape categories in Figure 8. We also plot the sorted number of occurrences of each part (bottom) in both semantic levels. In the coarse-grained semantic level, parts occurrences are concentrated on a smaller number of parts, while the distribution for the fine-grained level is clearly long-tailed. The average number of parts per object is also equalized across shape categories in the coarse-grained level, while some shape categories present a significantly higher number of parts per object in the fine-grained level like cars and bicycles. Overall, the coarse-grained semantic level provides a more balanced distribution of parts across shape categories, while the fine-grained semantic level provides a more detailed description of the composition of shapes.
@@ III-D Rendering
All images are rendered in a 256x256 resolution, and 2D images are encoded in the PNG format.
Depth maps are stored in the OpenEXR format to accommodate absolute distances to the image plane, which are represented using floating-point values.
Viewpoints. Each stylized shape is rendered from multiple perspectives: 4 canonical viewpoints and 4 random viewpoints. We first translate each shape above the $z=0$ plane. Camera viewpoints are defined in spherical coordinates $(\phi,\theta)$ where the origin is set to the center of the shape’s bounding box, which we note $\mathbf{o}_{c}$. The camera is rotated around $\mathbf{o}_{c}$ by $\phi$ and $\theta$. Canonical viewpoints are distributed evenly around the shape with a fixed elevation $\theta$. We set the base spherical angle $\phi$ to 40 degrees and then increment it by 90 degrees for each of the four viewpoints, while keeping $\theta$ fixed at 0 degrees. Random viewpoints are sampled uniformly from an upper hemisphere above the plane. We randomly sample $\phi$ from the range $[0,2\pi]$ and $\theta$ from the range $[-\frac{1}{3}\pi,\frac{1}{3}\pi]$. Using the obtained $\phi$ and $\theta$ angles, we define the position and orientation of the camera. The camera’s initial position denoted as $\mathbf{c_{0}}$ is rotated around $\mathbf{o_{c}}$. The orientation of the camera is then adjusted to ensure that the image plane is centered on $\mathbf{o_{c}}$. Extrinsic and intrisic camera parameters are recorded for each view and are provided alongside the rendered images. The sampling procedure of camera parameters is detailed in Algorithm 1.
III-E Toolbox
@@
Algorithm 1  Sampling camera parameters
Input:
  • $\mathbf{obj}\in\mathbb{R}^{N\times 3}$: Object to render.
  • $\mathbf{c_{0}}\in\mathbb{R}^{3}$: Base position of the camera.
  • $v_{k}\in\llbracket 0,3\rrbracket$: View identifier.
[…]
Output:
  • $\mathbf{m_{w}}\in\mathbb{R}^{4\times 4}$: Camera transformation matrix.
[…]
2.     $\mathbf{o_{c}}\leftarrow\text{bounding\_box\_center}(\mathbf{obj})$
3.     if is_random_view then
4.         $\mathbf{\theta}\sim\mathcal{U}{\left[-\frac{1}{3}\pi,\frac{1}{3}\pi\right]},\ \mathbf{\phi}\sim\mathcal{U}{\left[0,2\pi\right]}$
5.     else
6.         $\mathbf{\theta}\leftarrow 0,\ \mathbf{\phi}\leftarrow\frac{2}{9}\pi+\frac{\pi}{2}v_{k}$
7.     $\mathbf{c_{1}}\leftarrow\text{rotate\_point}(\mathbf{c_{0}};\ \mathbf{o_{c}},\mathbf{\phi},\mathbf{\theta})$
8.     $\mathbf{m_{w}}\leftarrow\text{look\_at}(\mathbf{c_{1}},\mathbf{o_{c}})$
9.     return $\mathbf{m_{w}}$triangleright Camera transformation matrix
IV Experiments
IV-A Classification and Segmentation
@@
Part segmentation. We conduct 3D part segmentation experiments on pointclouds and 2D renders to assess the difficulty of this task on our dataset. We provide results for both fine-grained and coarse-grained 3D part segmentation in Table III. We report pointwise accuracy (shape-agnostic) and mIOU for each model. For mIOU, we consider the shape-informed version of the metric where we restrict the set of predicted parts to the parts that are present in the ground-truth shape category, and the shape-agnostic version where all possible parts are considered. We also report results with and without using a shape prior during training and inference for PCT [28], PointNet++ [29] and CurveNet [27]. We note that getting accurate part segmentations without RGB information is challenging but remains possible. Without using a shape prior, CurveNet [27] reaches a shape-agnostic mIOU of 53.09% on fine-grained part segmentation.
In this setting, the model has to perform the challenging task of point-wise part classification from a set of 275 possible parts.
@@
Overall, a large gap exists (around 30 accuracy points across models) between shape-informed and shape-agnostic mIOU, highlighting the difficulty of the task over the full space of possible parts. The task of coarse-grained part segmentation is easier, as the model only has to perform part classification from a set of 43 possible parts. In this setting, CurveNet [27] reaches a shape-agnostic mIOU of 76.32%. For 2D fine and coarse part segmentation, we report results for SegFormer [30] in Table V, alongside material segmentation results. We obtain a mIOU of 52.24% for fine-grained part segmentation, 73.35% for coarse-grained part segmentation, and 82.45% for material segmentation.
IV-B Grounded Compositional Recognition (GCR)
  • Value. Proportion of correctly predicted part-material pairs.
We consider a part to be correctly segmented if the predicted part segmentation mask has an intersection over union (IoU) of at least 0.5 with the ground-truth part segmentation mask.
Formally, for a test set with $N$ shapes, let $y_{i}$ and $\hat{y}_{i}$ denote the true and predicted shape categories for shape $i$, respectively. Additionally, let $\mathcal{C}_{i}=\{(p_{1},m_{1}),(p_{2},m_{2}),\ldots,(p_{K_{i}},m_{K_{i}})\}$ represent the set of true part-material pairs for shape $i$, and $\hat{\mathcal{C}}_{i}$ the corresponding predicted pairs. Let $\mathcal{P}_{i}$ denote the set of all parts in shape $i$.
Using these definitions, we define the metrics as follows.
Shape Accuracy measures the proportion of correctly identified shape categories:
Shape $\textsc{Shape Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}(y_{i}=\hat{y}_{i})$
Value computes the average proportion of correctly predicted part-material pairs per shape:
$\textsc{V}=\frac{1}{N}\sum_{i=1}^{N}\frac{|\mathcal{C}_{i}\cap\hat{\mathcal{C}}_{i}|}{|\mathcal{C}_{i}|}$
Value-all requires perfect prediction of all part-material pairs:
$\textsc{Vall}=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}(\mathcal{C}_{i}=\hat{\mathcal{C}}_{i})$
With $\mathds{1}(\cdot)$ being the indicator function. Note that performing well on these metrics only requires part-material labels to be predicted, and does not require part segmentation masks. For the grounded variants, we define $\mathcal{S}_{i}$ as the set of correctly segmented parts in shape $i$, that is:
$S_{i}=\{p\in\mathcal{P}_{i}:\text{IoU}(p,\hat{p})\geq 0.5\}$
The $\text{IoU}(p,\hat{p})$ function computes the intersection over union between the predicted part $\hat{p}$ and the ground-truth part $p$. Depending on the modality, the $\text{IoU}(\cdot)$ function is defined either as the pixelwise IoU for 2D images or the pointwise IoU for 3D pointclouds. This allows us to define grounded variants of the metrics as follows:
Grounded-value which measures correctly predicted part-material pairs that are also correctly segmented:
$\textsc{GV}=\frac{1}{N}\sum_{i=1}^{N}\frac{|\{(p,m)\in\mathcal{C}_{i}:p\in\mathcal{S}_{i}\land(p,m)\in\hat{\mathcal{C}}_{i}\}|}{|\mathcal{C}_{i}|}$
Grounded-value-all which requires all part-material pairs to be correctly predicted and all parts to be correctly segmented:
$\textsc{GVall}=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}\left((\mathcal{C}_{i}=\hat{\mathcal{C}}_{i})\land(\forall p\in\mathcal{P}_{i}:p\in S_{i})\right)$
Note that Value and Grounded-value are both evaluated at the shape level: we divide the number of correctly identified (resp. grounded) part-material pairs by the total number of parts appearing in each shape, and then average across all samples. Value-all is thus upper bounded by Value, and Grounded-value-all by Grounded-value.